Org Pvt. Ltd. is considering two mutually exclusive capital investments. The project’s expected net cash flows are as follows:

Expected Cash Flows:

Year Project A Project B

0 -400 -575

1 95 150

2 110 200

3 118 250

4 125 275

5 140 230

6 150 180


a. If you were told that each project’s cost of capital was 10%, which project should be selected using the NPV criteria?

b. What is each project’s IRR?

c. What is the regular payback period for these two projects?

d. What is the profitability index for each project if the cost of capital is 12%?


1
Expert's answer
2021-05-12T16:56:30-0400

(a)


Since NPV is more of Project B as compared to Project A, hence it is

 advisable to accept Project B


(b)



(c)Calculating the regular payback period for Project A:



 payback period"=A+\frac{B}{C}=3+\frac{77}{125}=3.616 \space years"

Thus, the payback period for Project A is 3.616 years or 3.62 years (rounded off).

 

Calculating the regular payback period for Project B:


"=A+\frac{B}{C}=2+\frac{225}{250}=2.0909 \space years"

Thus, the payback period for Project B is 2.0909 years or 2.09 years. 


  • The variable A denotes the last year with a negative cumulative cash flow;
  • The variable B denotes the value without negative symbol cumulative net cash flow at the end of year A; and
  • The variable C denotes the gross cash flow during the year following year A


(d)


where,

PV Factor=1/(1+Cost of capital)^number of years

Discounted cash flow=Cash flow*PV factor

Present Value of total cash inflows=Sum of all the cash inflows


Profitability Index"=\frac{Present\space Value\space of\space cash \space in\space flows}
{Present \space Value\space of\space cash\space outflows}"


"Project A =\frac{491.377}{400}=1.23"

"Project B =\frac{867.782}{575}=1.51"